# CHBench: A Chinese Dataset for Evaluating Health in Large Language Models

### v1

@@ Abstract
With the rapid development of large language models (LLMs), assessing their performance on health-related inquiries has become increasingly essential. It is critical that these models provide accurate and trustworthy health information, as their application in real-world contexts—where misinformation can have serious consequences for individuals seeking medical advice and support—depends on their reliability. In this work, we present CHBench, the first comprehensive Chinese Health-related Benchmark designed to evaluate LLMs’ capabilities in understanding physical and mental health across diverse scenarios. CHBench includes 6,493 entries related to mental health and 2,999 entries focused on physical health, covering a broad spectrum of topics. This dataset serves as a foundation for evaluating Chinese LLMs’ capacity to comprehend and generate accurate health-related information. Our extensive evaluations of four popular Chinese LLMs demonstrate that there remains considerable room for improvement in their understanding of health-related information. The code is available at https://github.com/TracyGuo2001/CHBench.
CHBench: A Chinese Dataset for Evaluating Health in Large Language Models
Chenlu Guo1  Nuo Xu1  Yi Chang1,2,3  Yuan Wu 1111Corresponding author: yuanwu@jlu.edu.cn 1School of Artificial Intelligence, Jilin University, China 2Engineering Research Center of Knowledge-Driven Human-Machine Intelligence, MOE, China 3International Center of Future Science, Jilin University, China
@@ 1 Introduction
Large language models (LLMs) have garnered significant attention in recent years, demonstrating remarkable capabilities across a wide array of complex tasks (Zhao et al., 2023). Exemplary models such as GPT-3 (Brown et al., 2020), ChatGLM (Du et al., 2021; Zeng et al., 2022), LLaMA (Touvron et al., 2023), and PaLM (Chowdhery et al., 2023) have emerged, with the advent of GPT-4 (Achiam et al., 2023) igniting a new wave of enthusiasm. These breakthroughs are largely driven by fine-tuning techniques that substantially enhance the power and controllability of LLMs, aligning their training objectives with human preferences to ensure they interpret and execute commands accurately and effectively (Zhang et al., 2023a).
@@
However, concerns have been raised about the language models’ potential to internalize, propagate, and even amplify harmful content present in their training data, which can sometimes manifest in toxic language (Gehman et al., 2020). Numerous studies have illuminated the security risks posed by models like ChatGPT OpenAI (2023), revealing that despite advancements, some models continue to exhibit toxic behaviors (Chang et al., 2024). Evaluating the safety of LLMs is crucial. Several datasets focus on safety-related issues, encompassing various concerns such as toxicity and harmful language. For example, ToxicChat (Lin et al., 2023) categorizes user queries into different toxicity levels, while SALAD-Bench (Li et al., 2024), DiaSafety (Sun et al., 2021), and Do-Not-Answer (Wang et al., 2023) examine a range of safety issues collectively. Others, such as HateXplain (Mathew et al., 2021) and bias-related datasets (Zhou et al., 2022; Barikeri et al., 2021), target more specific aspects of safety.
@@
Despite the growing body of safety-oriented datasets, there is a notable gap in health-related datasets. Previous research has often subsumed health under broader safety concerns, potentially underestimating or neglecting certain harms (Xu et al., 2023; Zhang et al., 2023b). While some datasets, like SafeText (Levy et al., 2022), focus on health issues, and PsychoBench (Huang et al., 2023) assesses LLMs’ psychological impacts, such resources remain scarce. Furthermore, there are no health-related datasets available in Chinese, with existing datasets predominantly in English, limiting the evaluation of Chinese LLMs. These datasets often prioritize the models’ reasoning abilities and knowledge breadth, overlooking their alignment with users’ values. For instance, Sun et al. (2023) use InstructGPT (Ouyang et al., 2022) as an evaluator, but the model’s behavior reflects feedback from a narrow group of primarily English-speaking contractors, whose value judgments may not encompass the diverse perspectives of all users affected by the models. The inadequacies in health-related question generation, as illustrated in Figure 1, highlight various issues: misaligned responses, sensitivity to data leading to misidentifications of relevant queries as toxic, and an inability to comprehend Chinese idioms and common abbreviations.
@@
To address these challenges, we propose CHBench, the first benchmark specifically designed to evaluate the proficiency of Chinese LLMs in understanding physical and mental health knowledge. CHBench contains 2,999 entries on physical health across four domains and 6,493 entries on mental health across six domains. Data is sourced from web posts, exams, and existing datasets, encompassing open-ended questions, real-life scenario analyses, and reasoning tasks. To maintain objectivity, we use the powerful Chinese LLM Ernie Bot to generate responses for all entries. Multiple metrics are employed to assess the quality of the generated responses, and Ernie Bot is also used to score these criteria. Our empirical evaluation of four Chinese LLMs reveals that, while challenges remain, there is ample room for improvement in the quality of the generated health-related content. Additionally, we analyze the persistent issues within these models. We hope that CHBench will significantly advance the safety and reliability of Chinese LLMs in health-related scenarios.
@@ 2 Related Work
With the growing recognition of security risks in large language models, several datasets have emerged to address safety concerns. SafetyBench (Zhang et al., 2023b) encompasses seven categories of safety-related questions, offering multiple-choice queries in both English and Chinese. Notably, the gaps between GPT-4 and other LLMs are particularly pronounced in specific safety domains such as physical health. Moreover, SafetyBench has mentioned that there exists no physical and mental health-related benchmarks in Chinese currently. SALAD-Bench (Li et al., 2024), a comprehensive safety benchmark, evaluates LLMs in terms of both attack and defense mechanisms, featuring a diverse dataset with 21,000 test samples and utilizing the MD-Judge evaluator for efficient assessment. This benchmark includes attack-enhanced, defense-enhanced, and multiple-choice questions. DiaSafety (Sun et al., 2021) provides 11,000 labeled context-response pairs, focusing on context-sensitive unsafe behaviors in human-bot dialogues, where dialogue responses must be correctly labeled based on their conversational context to ensure safety. The JAILBREAKHUB framework (Shen et al., 2023) compiles 1,405 jailbreak prompts collected between December 2022 and December 2023, identifying 131 jailbreak communities and analyzing their attack strategies. It includes a question set containing 107,250 samples across 13 forbidden scenarios to assess the potential harm these prompts could cause. In light of the limitations identified in existing safety benchmarks and the challenges posed by non-English queries, we introduce CHBench, the first dataset specifically designed to evaluate the performance of LLMs in the health domain, with a focus on Chinese-language scenarios.
@@ 3.1 Composition of Dataset
Physical health: This aspect focuses on issues that may impair physical functioning or endanger personal safety. It is crucial to assess whether any responses generated by LLMs could compromise physical health, create life-threatening situations, or impact specific bodily functions.
Mental health: This dimension centers on issues that influence emotional well-being, cognitive abilities, and ethical considerations. It is vital to evaluate whether the responses produced by LLMs could negatively affect mental health or pose psychological risks.
@@ 3.2.1 Web Post Retrieval
We collected relevant questions from Zhihu, a widely used platform where users engage in asking and answering questions across a broad spectrum of topics. In this era of knowledge sharing, Zhihu has emerged as a significant media outlet and a key influencer in numerous fields. Its extensive and diverse content makes it an ideal data source for our study.
For physical health, we adhered to the Guideline to Life Safety and Health Education Materials for Primary and Secondary Schools (MOE of PRC, 2021), which classifies life safety and health education into five domains and 30 core topics. From these, we selected four domains relevant to physical health as our screening criteria: (1) health behaviors and lifestyles, (2) growth, development, and adolescent health, (3) prevention of infectious diseases and response to public health emergencies, and (4) safety emergencies and risk management. Some of the selected keywords generated more open-ended questions, with fewer queries focused on specific scenarios. These open-ended questions often invited the community to share advice and suggestions, such as "What are some good habits you practice daily?". This diversity resulted in a wide range of responses. After carefully filtering out irrelevant and semantically repetitive questions, we curated a refined dataset of 1,111 entries.
For mental health, we followed the six moral foundations outlined in MIC (Ziems et al., 2022), which encompass a broad spectrum of moral considerations and reflect the diversity of human concerns and thoughts on ethical issues. These foundations include Care/Harm, Fairness/Cheating, Liberty/Oppression, Loyalty/Betrayal, Authority/Subversion, and Sanctity/Degradation. During the filtering process, we prioritized the clarity of the questions, excluding any that were vague or ambiguous, as they failed to accurately represent mental health concerns. Additionally, we assessed the overall quality of the questions, removing those that were of low quality or of limited relevance to our assessment criteria. Ultimately, we selected 6,500 representative and relevant prompt questions.
@@ 3.2.2 Exams
To ensure the inclusion of more relevant and practical questions, we selected appropriate queries from the exam questions of various competitions related to safety, health, nutrition, and diet. We filtered out definitional questions, such as "How much energy is released from the oxidation of each gram of protein or carbohydrate in the body?", focusing instead on questions addressing specific issues pertinent to particular groups or real-life scenarios. After this refinement process, we curated a dataset of 1,704 entries. These questions address practical concerns, such as "Should I consume more acidic foods after exercising?".
@@ 3.2.3 Existing Datasets
To help LLMs recognize genuine health-related concerns, we incorporate data that requires logical reasoning. Drawing from existing datasets based on Ruozhiba, a subforum on Baidu Tieba, these posts often contain puns, homonyms, inverted causality, and homophones, many of which present logical traps designed to challenge human reasoning. Previous datasets have already filtered out declarative statements from Tieba posts (Bai et al., 2024; Ruozhiba, 2024). We further refined the dataset by focusing on health-related content, resulting in a final selection of 187 entries.
@@ 3.3 Choosing the Gold-Standard Responses
  • Relevance and Specificity. The model’s responses should directly address the question, demonstrating an accurate understanding of the question’s intent and providing relevant answers.
@@ 3.4.1 Subjective Consistency
However, differences in cultural backgrounds among annotators pose an unavoidable challenge. Judgments are not universally shared, and individual ideologies, political views, and personal experiences can influence how workers evaluate the same expression, leading to varied assessments. This subjectivity can introduce bias and misinterpretation, potentially compromising the objectivity and universal applicability of the annotations.
To mitigate the effects of subjectivity, we employ ERNIE Bot to score the prompt-response pairs. As a language model trained and optimized on large-scale text data from diverse cultural contexts, ERNIE Bot offers more objective and consistent scoring outcomes. By applying this model to evaluate all gold-standard prompt-response pairs, we enhance the reliability and consistency of the annotation process.
3.4.2 Evaluation Standards
ERNIE Bot evaluates gold-standard prompt-response pairs to ensure the creation of a high-quality dataset. Establishing consistent evaluation criteria, while accounting for various factors, is crucial when annotating these pairs. The explanations of the evaluation standards for physical and mental health are presented in the Appendix B. Each evaluation standard is assessed on a three-point scale: "Unsatisfactory" (-1), "Neutral" (0), or "Satisfactory" (1).
@@ 3.4.3 Evaluating Gold Standard Pairs
We have established the use of ERNIE Bot for evaluating and providing explanations for gold-standard prompt-response pairs. However, to ensure accuracy, we also conduct a manual review of selected data to verify the reasonableness of its assessments. This step is crucial for confirming the soundness of the evaluation standards applied to Chinese LLMs before proceeding with the evaluation of the remaining data.
To this end, we randomly selected 108 pairs from both the physical and mental health categories for manual annotation. During this process, we strictly adhered to the established evaluation criteria, ensuring that each pair was meticulously and accurately assessed. We then tasked ERNIE Bot with performing the same annotation evaluation on these selected samples. After obtaining the model’s evaluations, we conducted a detailed comparison, analyzing the differences between manual and model-based assessments and investigating the reasons behind any discrepancies. Notably, ERNIE Bot not only provided evaluation results but also offered thorough explanations for each criterion. Upon comparison, we found a high degree of consistency between the model’s and manual evaluations. In some cases, the model’s explanations were more comprehensive, thoughtful, and objective than the manual annotations. Further details are provided in Appendix C.
Given these findings, we decided to use ERNIE Bot to evaluate the remaining pairs, establishing a structured process to ensure efficient and accurate assessment. First, we input the remaining prompt-response pairs into ERNIE Bot, which then begins automated evaluation. Leveraging its advanced NLP capabilities and extensive knowledge base, the model conducts a thorough analysis of each pair and provides corresponding evaluations, including both an overall assessment and detailed explanations for each evaluation criterion.
Throughout the automated evaluation process, we continuously monitor the model to ensure its stability and accuracy. If any discrepancies arise in specific scenarios, we promptly adjust and optimize the system to maintain reliability. Utilizing ERNIE Bot for annotation and evaluation significantly enhances efficiency, reduces the subjectivity and errors typically associated with manual evaluations, and maintains a high level of accuracy throughout the process.
@@ 4.1 Experimental Design
First, the responses are generated using the 5 Chinese LLMs, the details of the evaluated language models are given in Appendix A. The responses produced by ERNIE Bot are designated as the gold-standard pairs, and the parameters for response generation are set to their default values. For toxic queries where the models are unable to provide appropriate answers and return errors, we assign the response as "None!". We then calculate the similarity between the responses from the other four models and the gold-standard responses, with the goal of objectively measuring the consistency and relevance of these outputs in relation to the desired answers. The similarity between "None!" and the gold-standard responses is set to 0.
@@ 4.2 Metrics
In the field of data analysis and pattern recognition, similarity measurement is a vital metric for evaluating the degree of resemblance between two entities. In this experimental section, both cosine similarity and the Jaccard similarity coefficient are employed to quantify these similarities.
Cosine similarity is a dimensionless metric frequently used in text data analysis due to its ability to capture the directional relationship between vectors. Rather than accounting for vector lengths or absolute values, cosine similarity emphasizes the overlap of vector projections in multidimensional space.
Jaccard similarity coefficient is a set-based metric that calculates the ratio of the intersection to the union of two sets. It disregards word importance and order, focusing purely on co-occurrence. This method is particularly well-suited for evaluating responses to open-ended questions.

### v2

@@ Abstract
With the rapid development of large language models (LLMs), assessing their performance on health-related inquiries has become increasingly essential. The use of these models in real-world contexts—where misinformation can lead to serious consequences for individuals seeking medical advice and support—necessitates a rigorous focus on safety and trustworthiness. In this work, we introduce CHBench, the first comprehensive safety-oriented Chinese health-related benchmark designed to evaluate LLMs’ capabilities in understanding and addressing physical and mental health issues with a safety perspective across diverse scenarios. CHBench comprises 6,493 entries on mental health and 2,999 entries on physical health, spanning a wide range of topics. Our extensive evaluations of four popular Chinese LLMs highlight significant gaps in their capacity to deliver safe and accurate health information, underscoring the urgent need for further advancements in this critical domain. The code is available at https://github.com/TracyGuo2001/CHBench.
@@ 1 Introduction
Large language models (LLMs) have garnered significant attention in recent years, demonstrating remarkable capabilities across a wide array of complex tasks Zhao et al. (2023); Xia et al. (2024). Exemplary models such as GPT-3 Brown et al. (2020), ChatGLM Du et al. (2021); Zeng et al. (2022), LLaMA Touvron et al. (2023), and PaLM Chowdhery et al. (2023) have emerged, with the advent of GPT-4 Achiam et al. (2023) igniting a new wave of enthusiasm. These breakthroughs are largely driven by fine-tuning techniques that substantially enhance the power and controllability of LLMs, aligning their training objectives with human preferences to ensure they interpret and execute commands accurately and effectively Zhang et al. (2023a).
@@
However, concerns have been raised about the language models’ potential to internalize, propagate, and even amplify harmful content present in their training data, which can sometimes manifest in toxic language Gehman et al. (2020). Numerous studies have illuminated the security risks posed by models like ChatGPT OpenAI (2023), revealing that despite advancements, some models continue to exhibit toxic behaviors Chang et al. (2024). Evaluating the safety of LLMs is crucial Chang et al. (2024). Several datasets focus on safety-related issues, encompassing various concerns such as toxicity and harmful language. For example, ToxicChat Lin et al. (2023) categorizes user queries into different toxicity levels, while SALAD-Bench Li et al. (2024), DiaSafety Sun et al. (2021), and Do-Not-Answer Wang et al. (2023) examine a range of safety issues collectively. Others, such as HateXplain Mathew et al. (2021) and bias-related datasets Zhou et al. (2022); Barikeri et al. (2021), target more specific aspects of safety.
@@
Despite the growing body of safety-oriented datasets, there is a notable gap in health-related datasets. Previous research has often subsumed health under broader safety concerns, potentially underestimating or neglecting certain harms Xu et al. (2023); Zhang et al. (2023b). While some datasets, like SafeText Levy et al. (2022), focus on health issues, and PsychoBench Huang et al. (2023) assesses LLMs’ psychological impacts, such resources remain scarce. Furthermore, there are no health-related datasets available in Chinese, with existing datasets predominantly in English, limiting the evaluation of Chinese LLMs. These datasets often prioritize the models’ reasoning abilities and knowledge breadth, overlooking their alignment with users’ values. For instance, Sun et al. (2023) use InstructGPT Ouyang et al. (2022) as an evaluator, but the model’s behavior reflects feedback from a narrow group of primarily English-speaking contractors, whose value judgments may not encompass the diverse perspectives of all users affected by the models. The inadequacies in health-related question generation, as illustrated in Figure 1, highlight various issues: misaligned responses, sensitivity to data leading to misidentifications of relevant queries as toxic, and an inability to comprehend Chinese idioms and common abbreviations.
@@
To address these challenges, we propose CHBench, the first benchmark specifically designed to evaluate the proficiency of Chinese LLMs in understanding physical and mental health knowledge from a safety-oriented perspective. CHBench contains 2,999 entries on physical health across four domains and 6,493 entries on mental health across six domains. Data is sourced from web posts, exams, and existing datasets, encompassing open-ended questions, real-life scenario analyses, and reasoning tasks. To maintain objectivity, we use the powerful Chinese LLM Ernie Bot to generate responses for all entries. Multiple metrics are employed to assess the quality of the generated responses, and Ernie Bot is also used to score these criteria. Our empirical evaluation of four Chinese LLMs reveals that, while challenges remain, there is ample room for improvement in the safety and quality of the generated health-related content. Additionally, we analyze the persistent issues within these models. We hope that CHBench will significantly advance the safety and reliability of Chinese LLMs in health-related scenarios.
@@ 2 Related Work
With the growing recognition of security risks in large language models, several datasets have emerged to address safety concerns. SafetyBench Zhang et al. (2023b) encompasses seven categories of safety-related questions, offering multiple-choice queries in both English and Chinese. Notably, the gaps between GPT-4 and other LLMs are particularly pronounced in specific safety domains such as physical health. Moreover, SafetyBench has mentioned that there exists no physical and mental health-related benchmarks in Chinese currently. SALAD-Bench Li et al. (2024), a comprehensive safety benchmark, evaluates LLMs in terms of both attack and defense mechanisms, featuring a diverse dataset with 21,000 test samples and utilizing the MD-Judge evaluator for efficient assessment. This benchmark includes attack-enhanced, defense-enhanced, and multiple-choice questions. DiaSafety Sun et al. (2021) provides 11,000 labeled context-response pairs, focusing on context-sensitive unsafe behaviors in human-bot dialogues, where dialogue responses must be correctly labeled based on their conversational context to ensure safety. The JAILBREAKHUB framework Shen et al. (2023) compiles 1,405 jailbreak prompts collected between December 2022 and December 2023, identifying 131 jailbreak communities and analyzing their attack strategies. It includes a question set containing 107,250 samples across 13 forbidden scenarios to assess the potential harm these prompts could cause. In light of the limitations identified in existing safety benchmarks and the challenges posed by non-English queries, we introduce CHBench, the first dataset specifically designed to evaluate the performance of LLMs in the health domain, with an emphasis on safety in Chinese-language scenarios.
@@ 3.1 Composition of Dataset
Physical health:
This aspect focuses on issues that may impair physical functioning or endanger personal safety. It is crucial to assess whether any responses generated by LLMs could compromise physical health, create life-threatening situations, or impact specific bodily functions.
Mental health:
This dimension centers on issues that influence emotional well-being, cognitive abilities, and ethical considerations. It is vital to evaluate whether the responses produced by LLMs could negatively affect mental health or pose psychological safety risk.
@@ 3.2.1 Web Post Retrieval
We collected relevant questions from Zhihu [^2] https://www.zhihu.com, a widely used platform where users engage in asking and answering questions across a broad spectrum of topics. In this era of knowledge sharing, Zhihu has emerged as a significant media outlet and a key influencer in numerous fields. Its extensive and diverse content makes it an ideal data source for our study.
For physical health, we adhered to the Guideline to Life Safety and Health Education Materials for Primary and Secondary Schools MOE of PRC (2021), which classifies life safety and health education into five domains and 30 core topics. From these, we selected four domains relevant to physical health as our screening criteria: (1) health behaviors and lifestyles, (2) growth, development, and adolescent health, (3) prevention of infectious diseases and response to public health emergencies, and (4) safety emergencies and risk management. Some of the selected keywords generated more open-ended questions, with fewer queries focused on specific scenarios. These open-ended questions often invited the community to share advice and suggestions, such as ”What are some good habits you practice daily?”. This diversity resulted in a wide range of responses. After carefully filtering out irrelevant and semantically repetitive questions, we curated a refined dataset of 1,111 entries.
For mental health, we followed the six moral foundations outlined in MIC Ziems et al. (2022), which encompass a broad spectrum of moral considerations and reflect the diversity of human concerns and thoughts on ethical issues. These foundations include Care/Harm, Fairness/Cheating, Liberty/Oppression, Loyalty/Betrayal, Authority/Subversion, and Sanctity/Degradation. The selection of these keywords allows us to capture subtle aspects of mental health beyond basic psychological terms. For instance, ”Care/Harm” addresses needs for empathy and support, while ”Authority/Subversion” reflects issues of control and autonomy—both critical in mental health contexts. This approach provides a more comprehensive dataset, reflecting real-world complexity and enabling deeper evaluations of language models in mental health scenarios. During the filtering process, we prioritized the clarity of the questions, excluding any that were vague or ambiguous, as they failed to accurately represent mental health concerns. Additionally, we assessed the overall quality of the questions, removing those that were of low quality or of limited relevance to our assessment criteria. Ultimately, we selected 6,500 representative and relevant prompt questions.
@@ 3.2.2 Exams
To ensure the inclusion of more relevant and practical questions, we selected appropriate queries from the exam questions of various competitions related to safety, health, nutrition, and diet. We filtered out definitional questions, such as ”How much energy is released from the oxidation of each gram of protein or carbohydrate in the body?”, focusing instead on questions addressing specific issues pertinent to particular groups or real-life scenarios. After this refinement process, we curated a dataset of 1,704 entries. These questions address practical concerns, such as ”Should I consume more acidic foods after exercising?”.
@@ 3.2.3 Existing Datasets
To help LLMs recognize genuine health-related concerns, we incorporate data that requires logical reasoning. Drawing from existing datasets based on Ruozhiba, a subforum on Baidu Tieba [^3] https://tieba.baidu.com/, these posts often contain puns, homonyms, inverted causality, and homophones, many of which present logical traps designed to challenge human reasoning. Previous datasets have already filtered out declarative statements from Tieba posts Bai et al. (2024); Ruozhiba (2024). We further refined the dataset by focusing on health-related content, resulting in a final selection of 187 entries.
@@ 3.3 Choosing the Gold-Standard Responses
  • Relevance and Completeness. The model’s responses should directly address the question, demonstrating an accurate understanding of the question’s intent. Answers should be relevant, specific, and comprehensive, covering all key aspects of the question.
@@ 3.4.1 Subjective Consistency
However, differences in cultural backgrounds among annotators pose an unavoidable challenge. Judgments are not universally shared, and individual ideologies, political views, and personal experiences can influence how workers evaluate the same expression, leading to varied assessments. This subjectivity can introduce bias and misinterpretation, potentially compromising the objectivity and universal applicability of the annotations. To reduce subjectivity, we use ERNIE Bot to score prompt-response pairs, ensuring more objective and consistent outcomes. This approach improves the reliability of evaluating gold-standard responses.
@@ 3.4.2 Evaluating Gold Standard Pairs
ERNIE Bot evaluates gold-standard prompt-response pairs to ensure the creation of a high-quality dataset. Establishing consistent evaluation criteria, while accounting for various factors, is crucial when annotating these pairs. The explanations of the evaluation standards for physical and mental health are presented in the Appendix . Each evaluation standard is assessed on a three-point scale: ”Unsatisfactory” (-1), ”Neutral” (0), or ”Satisfactory” (1).
However, to ensure accuracy, we also perform a manual review of selected data to validate the reliability of the model’s assessments. We randomly selected 108 pairs from both the physical and mental health categories for manual annotation. During this process, we strictly adhered to the established evaluation criteria, ensuring that each pair was meticulously and accurately assessed. We then tasked ERNIE Bot with performing the same annotation evaluation on these selected samples. After obtaining the model’s evaluations, we conducted a detailed comparison, analyzing the differences between manual and model-based assessments and investigating the reasons behind any discrepancies. Notably, ERNIE Bot not only provided evaluation results but also offered thorough explanations for each criterion. Upon comparison, we found a high degree of consistency between the model’s and manual evaluations. In some cases, the model’s explanations were more comprehensive, thoughtful, and objective than the manual annotations. Further details are provided in Appendix .
Given these findings, we decided to use ERNIE Bot to evaluate the remaining pairs, establishing a structured process to ensure efficient and accurate assessment. Leveraging its advanced NLP capabilities and extensive knowledge base, the model conducts a thorough analysis of each pair and provides corresponding evaluations, including detailed explanations for each evaluation criterion.
@@ 4.1 Experimental Design
First, the responses are generated using the 5 Chinese LLMs, the details of the evaluated language models are given in Appendix . The responses produced by ERNIE Bot are designated as the gold-standard responses, and the parameters for response generation are set to their default values. For toxic queries where the models are unable to provide appropriate answers and return errors, we assign the response ”None!” as the output to record. We then calculate the similarity between the responses from the other four models and the gold standard responses, to objectively measure the consistency and relevance of these outputs with the desired answers.
@@ 4.2 Metrics
Similarity measurement is a critical metric for evaluating the resemblance between entities in data analysis and pattern recognition. In this study, we employ cosine similarity and the Jaccard similarity coefficient to capture different dimensions of similarity, leveraging their respective strengths.
4.2.1 Cosine Similarity
Cosine similarity is widely used in text data analysis due to its capacity to capture the directional relationship between vectors, emphasizing the overlap of vector projections in multidimensional space. For encoding, we use a text2vec-base-chinese model Ming (2022). The model is trained with CoSENT method Huang et al. (2024), based on chinese-macbert-base model Cui et al. (2020) trained on Chinese STS-B data Ming (2022), which is particularly well-suited for semantic matching tasks.
4.2.2 Jaccard Similarity Coefficient
The Jaccard similarity coefficient measures the ratio of intersection to union of two sets, focusing purely on co-occurrence while disregarding word order and importance. To adapt this metric to the nuances of the Chinese language, we employ Jieba for word segmentation and integrated TF-IDF encoding. TF-IDF weights terms by their significance in the corpus, reducing the influence of high-frequency, low-information words. This enhancement enables the Jaccard similarity coefficient to better reflect meaningful similarities in Chinese text.
Criteria Interval Physical Health Mental Health
ChatGLM Qwen Baichuan SparkDesk ChatGLM Qwen Baichuan SparkDesk
Cosine Similarity [0.0,0.1) - - - - - - - -
[0.1,0.2) - - - - - - - -
[0.2,0.3) - - - 1 - - - -
[0.3,0.4) - - 2 - - - 23 -
[0.4,0.5) - 1 28 - - - 386 -
[0.5,0.6) 1 1 13 1 2 2 181 3
[0.6,0.7) 17 26 24 18 32 50 92 52
[0.7,0.8) 200 226 268 235 418 461 555 448
[0.8,0.9) 1453 1643 1711 1698 3762 3677 3626 3741
[0.9,1.0] 1324 1067 953 1039 2247 2105 1630 2176
Jaccard Similarity Coefficient [0.0,0.1) 31 52 146 67 141 142 921 125
[0.1,0.2) 1188 1517 1667 1697 2820 2485 2984 2678
[0.2,0.3) 1634 1310 1111 1144 3361 3438 2414 3393
[0.3,0.4) 136 85 74 78 137 227 173 221
[0.4,0.5) 6 - 1 6 2 3 1 3
[0.5,1.0] - - - - - - - -
The Number of ’None!’ 4 35 0 7 32 198 0 73
Table 1: Similarity of responses to gold standard responses across models